OpenStudy (anonymous):

ophia invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of her investment after x years: f(x) = 500(1.05)x What was the average rate of change of the value of Sophia's investment from the second year to the fourth year? 14.13 dollars per year 28.25 dollars per year 50.00 dollars per year 56.50 dollars per year

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi :)

OpenStudy (misty1212):

average is \[\frac{f(4)-f(2)}{4-2}\]

OpenStudy (anonymous):

i would use a calculator

OpenStudy (misty1212):

i would use wolfram, it is right on the computer

OpenStudy (anonymous):

i need an explaination on how to do this :(

OpenStudy (anonymous):

ill just yolo it, thanks though

jimthompson5910 (jim_thompson5910):

f(x) = 500(1.05)^x f(2) = 500(1.05)^2 ... replace every x with 2 now use a calculator to evaluate `500(1.05)^2` if you don't have one, you can use a free online one like this http://web2.0calc.com/ type in `500(1.05)^2` then hit enter or the equal sign. Tell me what you get so I can check to see if it's correct or not.

OpenStudy (anonymous):

thank you :)